# The Rate-Distortion-Perception-Classification Tradeoff: Joint Source Coding and Modulation via Inverse-Domain GANs

### v1

@@ Abstract
The joint source coding and modulation (JSCM) framework was enabled by recent developments in deep learning, which allows to automatically learn from data, and in an end-to-end fashion, the best compression codes and modulation schemes. In this paper, we show the existence of a strict tradeoff between channel rate, distortion, perception, and classification accuracy in a JSCM scenario. We then propose two image compression methods to navigate that tradeoff: an inverse-domain generative adversarial network (ID-GAN), which achieves extreme compression, and a simpler, heuristic method that reveals insights about the performance of ID-GAN. Experiment results not only corroborate the theoretical findings, but also demonstrate that the proposed ID-GAN algorithm significantly improves system performance compared to traditional separation-based methods and recent deep JSCM architectures.
@@
Image compression, joint source coding and modulation, generative adversarial networks, rate-distortion-perception-classification tradeoff.
@@ I Introduction
Traditional communication systems follow the celebrated source-channel coding theorem by Shannon [1], which states that source coding and channel coding can be designed separately without loss of optimality. Source coding removes redundant information from a signal, for example, by representing it in a different domain and zeroing out small coefficients. Channel coding, on the other hand, adds to the resulting compressed signal additional information, error-correcting codes, to make its transmission via a noisy channel more robust. Such a modular design, while optimal for memoryless ergodic channels with codes of infinite block length, becomes unsuitable for extreme scenarios, e.g., when bandwidth is highly limited or the channel varies rapidly. An example is underwater acoustic communication, in which multipath interference and noise are so large that the performance of separate source-channel coding schemes sharply drops below a certain signal-to-noise ratio (SNR), a phenomenon known as the cliff effect [2]. Traditional techniques like fast adaptive modulation and coding rarely work in such an environment, especially for long source bit sequences like images.
Joint source coding and modulation. The above problem can be addressed by designing the source coding and modulation schemes jointly, a framework known as joint source coding and modulation (JSCM). Enabled by deep learning, JSCM leverages training data to automatically learn the best compression codes and modulation schemes in an end-to-end fashion. Signals are thus directly mapped onto a complex-valued constellation space, and the resulting representation is transmitted through the noisy channel.
In the context of complex signals (like images), and under extreme compression requirements (as in underwater communication), the selection of the features to be compressed strikes different tradeoffs between different metrics: for example, optimizing for image reconstruction may reduce the perceptual quality of the reconstructed image or decrease the accuracy of a subsequent image classification algorithm. The study of such type of tradeoffs started with the seminal work in [3], which modified the rate-distortion framework [1] to study the tradeoff between perception and distortion metrics of image restoration algorithms. In this paper, we extend the study of such tradeoffs to a JSCM scenario. This requires considering not only vector signals, and thus the possibility to reduce their dimension, but also various metrics, including rate, distortion, perception, and classification performance.
The RDPC function. We introduce the rate-distortion-perception-classification (RDPC) function in a JSCM scenario. To define it, we consider an $n$-dimensional source signal $\bm{X}\in\mathbb{R}^{n}$ that can be drawn from one of $L$ classes:
@@
where $H_{l}$ represents the hypothesis that $\bm{X}$ is drawn from class $l$, which occurs with probability $p_{l}\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}(H_{l})$. The communication process is modeled as a Markov chain
[…]
where $\bm{X}$, $\widehat{\bm{X}}\in\mathbb{R}^{n}$ represent the source and reconstructed signals, and $\bm{Y}$, $\widehat{\bm{Y}}\in\mathbb{R}^{m}$, with $m<n$, represent the transmitted and received signals. The distribution $p_{\bm{Y}|\bm{X}}$ (resp. $p_{\widehat{\bm{Y}}|\bm{Y}}$ and $p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}$) characterizes the encoder (resp. channel and decoder). We assume the channel adds zero-mean Gaussian noise to $\bm{Y}$, i.e., $p_{\widehat{\bm{Y}}|\bm{Y}}(\widehat{\bm{y}}|\bm{y})=\mathcal{N}\big{(}\bm{y}\,,\bm{\Sigma}\big{)}$, where $\bm{\Sigma}$ is an $m\times m$ diagonal matrix whose $i$th diagonal entry $\bm{\Sigma}_{ii}>0$ represents the noise power of channel $i$.
@@
$R(D,P,C)=\begin{array}[t]{cl}\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\min}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq D\\[2.84544pt] &d(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq P\\[2.84544pt] &\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\leq C\,.\end{array}$ (7)
@@
Our approach. As existing characterizations of tradeoffs between, for example, distortion and perception [3], rate, distortion, and perception [4], or classification error, distortion, and perception [5], we show the existence of a tradeoff between rate, distortion, perception, and classification error. Our setup [cf. (2)] is more general than the ones in [3, 4, 5], as we consider vector signals (not necessarily scalar) and their compression in terms of dimensionality. Furthermore, channel capacity is evaluated at test time by injecting different levels of channel noise, obtaining different rates. We also establish a strict tradeoff between all the above quantities, i.e., that the function $R(D,P,C)$ is strictly convex in $D$, $P$, and $C$.
It is difficult to solve (7) in full generality. So, leveraging the capacity of generative adversarial networks (GANs) to model probability distributions [6], we propose to use inverse-domain GAN (ID-GAN) [7] to design an image compression algorithm that achieves both extremely high compression rates and good quality in terms of reconstruction, perception, and classification. Compared to the original GAN [6], ID-GAN [7] learns how to map not only a latent code to an image, but also an image to a latent code. To understand the behavior of this algorithm, we also propose a method that attempts to directly solve (7) under simple assumptions, e.g., that there are only two classes ($L=2$) and that the encoder and decoder are linear maps. Despite these simplifications, experimental results show the same type of behavior for both algorithms.
Contributions. We summarize our contributions as follows:
@@
  • We propose two algorithms to solve the JSCM tradeoff problem: one based on inverse-domain GAN (ID-GAN) [7], and a simpler one to understand its behavior. The proposed inverse-domain GAN algorithm can transmit images under extreme compression rates, handling low-capacity channels and preserving semantic information, perception quality, and reconstruction fidelity.
  • We upper bound the optimal value in (7) when the input signal is a Gaussian mixture model (GMM). To achieve this, we obtain a new bound on the Wasserstein-$1$ distance between GMMs in terms of their parameters.
  • Experiment results show that the proposed ID-GAN algorithm outperforms both a traditional method with source coding and modulation designed separately (JPEG+LDPC+BPSK) and recent deep JSCM algorithms like D-JSCC [2] and AE+GAN [8].
Organization. We overview related work in Section II and characterize the tradeoff problem (7) in Section III. Section IV then describes the proposed JSCM algorithms, whose performance is assessed in Section V. We conclude in Section VI.
@@ II Related Work
II-A Joint Source-Channel Coding and Modulation (JSCM)
As mentioned, JSCM schemes achieve better reconstruction than the classical source-channel separation method. They can be divided into two categories: methods that focus on channel transmission, and methods that focus on source coding.
JSCM for channel transmission. The main goal of JSCM for channel transmission is to mitigate the effects of channel noise. For example, [2] proposed a deep joint source-channel coding (D-JSCC) algorithm based on an autoencoder and showed that besides outputting images with quality superior to separation-based schemes, the algorithm exhibits graceful performance degradation in low SNR. [^3] Although the method in [2] is named JSCC, it maps the compressed images directly onto the constellation space (IQ domain). So it performs modulation rather than source-coding, and is therefore an JSCM method.
Inspired by D-JSCC, subsequent work has designed retrieval-oriented image compression schemes [9], used channel feedback to improve the quality of transmission [10], and considered an adaptive bandwidth to transmit information progressively [11]. Furthermore, [12] designed an end-to-end approach for D-JSCC with channel state information (CSI) feedback. The main idea is to apply a non-linear transform network to compress both the data and the CSI. Finally, [13] designed a scheme for orthogonal frequency division multiplexing (OFDM) transmission that directly maps the source images onto complex-valued baseband samples.
@@ II-A Joint Source-Channel Coding and Modulation (JSCM)
JSCM for source coding. Techniques used in JSCM for source coding vary according to the domain of the data, e.g., text, image, video, or multimodal data. For example, the JSCM system designed in [14] used a recurrent neural network for transmitting text. Also focusing on text transmission, [15] proposed a semantic communication system (DeepSC) based on a transformer and, to evaluate performance, also a novel metric to measure sentence similarity. DeepCS was extended in [16] for speech transmission. JSCM has also been applied to the transmission of multimodal data. For instance, [17] proposed a cooperative scheme to transmit audio, video, and sensor data from multiple end devices to a central server. And concentrating on text and images, [18] designed a coarse-to-fine multitask semantic model using an attention mechanism.
@@
GAN-based compression. Most algorithms for image transmission are based on autoencoders [19], e.g., [2, 9, 11, 10, 20]. Autoencoders, however, compress signals only up to moderate compression ratios. Although they achieve high-quality reconstruction, this is at the cost of communication efficiency. Extreme compression has been achieved instead by using generative adversarial networks (GANs) [6]. A GAN is a generative model that learns, without supervision, both a low-dimensional representation of the data and its distribution [21]. This gives them the potential to achieve extreme compression without undermining image perception quality. For example, [8] proposed an autoencoder-GAN (AE+GAN) image compression system in which the encoder and decoder are trained simultaneously. The resulting method can achieve extremely low bitrates. The algorithm we propose, ID-GAN, requires less training (as encoder and decoder are trained separately), but attains a performance similar or better than AE+GAN [8].
@@
Also related to our work, [22] proposed two algorithms, inverse-JSCC and generative-JSCC, to reconstruct images passed through a fixed channel with a high compression ratio. The inverse-JSCC algorithm views image reconstruction as an inverse problem and uses a powerful GAN model, StyleGAN-2 [23], as a regularizer together with a distortion loss that aligns with human perception, LPIPS [24]. It is thus an unsupervised method. Generative-JSCC transforms inverse-JSCC into a supervised method by learning the parameters of an encoder/decoder pair while keeping the parameters of StyleGAN-2 fixed. This work differs from ours in several ways. First, we consider not only distortion and perception metrics, but also classification accuracy and channel rate. In particular, the experiments in [22] do not consider any classification task. We also characterize the tradeoff between all these metrics. Second, our metric for perception, the Wasserstein-1 distance between the input and output distributions, differs from the LPIPS metric. Third, we train both the encoder and the decoder adversarially, while [22] uses a pre-trained GAN for the decoder. Finally, while training StyleGAN-2 in [22] (on a database of faces) requires tremendous computational resources, training our networks takes more modest resources.
II-B Tradeoff Analyses

### v2

@@ Abstract
The joint source-channel coding (JSCC) framework leverages deep learning to learn from data the best codes for source and channel coding. When the output signal, rather than being binary, is directly mapped onto the IQ domain (complex-valued), we call the resulting framework joint source coding and modulation (JSCM). We consider a JSCM scenario and show the existence of a strict tradeoff between channel rate, distortion, perception, and classification accuracy, a tradeoff that we name RDPC. We then propose two image compression methods to navigate that tradeoff: the RDPCO algorithm which, under simple assumptions, directly solves the optimization problem characterizing the tradeoff, and an algorithm based on an inverse-domain generative adversarial network (ID-GAN), which is more general and achieves extreme compression. Simulation results corroborate the theoretical findings, showing that both algorithms exhibit the RDPC tradeoff. They also demonstrate that the proposed ID-GAN algorithm effectively balances image distortion, perception, and classification accuracy, and significantly outperforms traditional separation-based methods and recent deep JSCM architectures in terms of one or more of these metrics.
@@
Image compression, joint source-channel coding, joint source coding and modulation, generative adversarial networks, rate-distortion-perception-classification tradeoff.
@@ I Introduction
Traditional communication systems follow the celebrated source-channel coding theorem by Shannon [cover1999elements], which states that source coding and channel coding can be designed separately without loss of optimality. Source coding removes redundant information from a signal, for example, by representing it in a different domain and zeroing out small coefficients. Channel coding, on the other hand, adds to the resulting compressed signal additional information, error-correcting codes, to make its transmission via a noisy channel more robust. Such a modular design, while optimal for memoryless ergodic channels with codes of infinite block length, becomes unsuitable for extreme scenarios, e.g., when bandwidth is highly limited or the channel varies rapidly. An example is underwater acoustic communication, in which multipath interference and noise are so large that the performance of separate source-channel coding schemes sharply drops below a certain signal-to-noise ratio (SNR), a phenomenon known as the cliff effect [deep-jscc]. Traditional techniques like fast adaptive modulation and channel coding rarely work in such an environment, especially for long source bit sequences like images.
Joint source coding and modulation. The above problem can be addressed by jointly designing the source coding, channel coding, and modulation schemes, a framework we call joint source coding and modulation (JSCM). JSCM directly maps signals to the IQ domain and generalizes joint source and channel coding (JSCC) [fresia2010joint], in which the output signal is binary rather than complex-valued.
In the context of large signals (like images), and under extreme compression requirements (as in underwater communication), the selection of the features to be compressed strikes tradeoffs between different metrics: for example, optimizing for image reconstruction may reduce the perceptual quality of the reconstructed image or decrease the accuracy of a subsequent image classification algorithm. The study of such type of tradeoffs started with the seminal work in [blau2018perception], which modified the rate-distortion framework [cover1999elements] to study the tradeoff between perception and distortion metrics of image restoration algorithms. In this paper, we extend the study of such tradeoffs to a JSCM scenario. This requires considering not only vector signals, and thus the possibility to reduce their dimension, but also various metrics, including rate, distortion, perception, and classification performance.
I-A The RDPC function and problem statement
We introduce the rate-distortion-perception-classification (RDPC) function in a JSCM scenario. To define it, we consider an $n$-dimensional source signal $\bm{X}\in\mathbb{R}^{n}$ that can be drawn from one of $L$ classes:
@@ I-A The RDPC function and problem statement
where $H_{l}$ represents the hypothesis that $\bm{X}$ is drawn from class $l$, which occurs with probability $p_{l}:=\mathbb{P}(H_{l})$. The communication process is modeled as a Markov chain
${H_{l}}$${\bm{X}}$${\bm{Y}}$${\widehat{\bm{Y}}}$${\widehat{\bm{X}}\,,}$$\scriptstyle{p_{\bm{X}|H_{l}}\,\,\,}$$\scriptstyle{p_{\bm{Y}|\bm{X}}}$$\scriptstyle{p_{\widehat{\bm{Y}}|\bm{Y}}}$$\scriptstyle{p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}}$ (2)
where $\bm{X}$, $\widehat{\bm{X}}\in\mathbb{R}^{n}$ represent the source and reconstructed signals, and $\bm{Y}$, $\widehat{\bm{Y}}\in\mathbb{R}^{m}$, with $m<n$, represent the transmitted and received signals. The distribution $p_{\bm{Y}|\bm{X}}$ (resp. $p_{\widehat{\bm{Y}}|\bm{Y}}$ and $p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}$) characterizes the encoder (resp. channel and decoder). We assume the channel adds zero-mean Gaussian noise to $\bm{Y}$, i.e., $p_{\widehat{\bm{Y}}|\bm{Y}}(\widehat{\bm{y}}|\bm{y})=\mathcal{N}\big{(}\bm{y}\,,\bm{\Sigma}\big{)}$, where $\bm{\Sigma}$ is an $m\times m$ diagonal matrix whose $i$th diagonal entry $\bm{\Sigma}_{ii}>0$ represents the noise power of channel $i$. Motivated by recent work on task-aware image compression [Hua22-SemanticsGuidedContrastiveJointSourceChannelCodingForImageTransmission, Lei23-ProgressiveDeepImageCompressionWithHybridContextsOfImageClassificationAndReconstruction, Hua24-ClassificationDrivenDiscreteNeuralRepresentationLearningForSemanticCommunications], the purpose of the communication channel in (2) is to transmit images that can be used across different tasks, each of which may have different requirements in terms of image fidelity, perception, or classification. Considering classification as a task justifies the assumption in (1) that $\bm{X}$ always belongs to a given class.
@@
$R(D,P,C)=\begin{array}[t]{cl}\underset{p_{\bm{Y}|\bm{X}},\,p_{\widehat{\bm{X}}|\widehat{\bm{Y}}},\,\bm{\Sigma}}{\min}&\sum_{i=1}^{m}\log\big{(}1+\frac{1}{\bm{\Sigma}_{ii}}\big{)}\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq D\\[2.84544pt] &d(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq P\\[2.84544pt] &\mathbb{E}\big{[}\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})\big{]}\leq C\,.\end{array}$ (7)
@@
I-B Our approach and contributions
As existing characterizations of tradeoffs between, for example, distortion and perception [blau2018perception], rate, distortion, and perception [RDPtradeoff], or classification error, distortion, and perception [CDP], we show the existence of a tradeoff between rate, distortion, perception, and classification error. Our setup [cf. (2)] is more general than the ones in [blau2018perception, RDPtradeoff, CDP], as we consider vector signals (not necessarily scalar) and their compression in terms of dimensionality. We also establish a strict tradeoff between all the above quantities, i.e., that the function $R(D,P,C)$ is strictly convex in $D$, $P$, and $C$.
It is difficult to solve (7) in full generality. So, under the assumption that the source signals are drawn from a Gaussian mixture model (GMM) with two classes ($L=2$) and that the encoder and decoder are linear maps, we design an algorithm, RDPCO, that directly attempts to solve (7). In addition, leveraging the capacity of generative adversarial networks (GANs) to model probability distributions [goodfellow2014generative], we also propose to use inverse-domain GAN (ID-GAN) [IndomainGAN] to design an image compression algorithm that achieves both extremely high compression rates and good quality in terms of reconstruction, perception, and classification. Compared to the original GAN [goodfellow2014generative], ID-GAN [IndomainGAN] learns how to map not only a latent code to an image, but also an image to a latent code. Despite several differences, experimental results show that algorithms RDPCO and ID-GAN exhibit a similar behavior. We summarize our contributions as follows:
@@ I-B Our approach and contributions
  • We propose two algorithms to solve the JSCM tradeoff problem: a simple algorithm (RDPCO) that directly solves the tradeoff problem but applies only under restrictive assumptions, and another based on inverse-domain GAN (ID-GAN) [IndomainGAN] which can transmit images under extreme compression rates, handling low-capacity channels and preserving semantic information, perception quality, and reconstruction fidelity. In particular, we port techniques from [IndomainGAN], originally applied to image editing, to a JSCM scenario.
  • We upper bound the optimal value in (7) when the input signal is a GMM and the encoder and decoder are linear. To achieve this, we derive a new bound on the Wasserstein-$1$ distance between GMMs in terms of their parameters. See Lemma 3.
  • Simulation results show that RDPCO and ID-GAN exhibit the same behavior and reveal further insights about the RDPC problem. In addition, the proposed ID-GAN algorithm achieves a better RDPC tradeoff than a traditional method with source coding and modulation designed separately (JPEG+LDPC+BPSK) and than AE+GAN [agustsson2019generative], a recent deep algorithm (modified to a JSCM scenario). It also achieves much better perception and classification accuracy than D-JSCC [deep-jscc], at the cost of a slight increase in distortion.
I-C Organization
We overview related work in Section II and characterize the tradeoff problem (7) in Section III. Section IV analyzes the RDPC tradeoff under GMM source signals and linear encoders/decoders, and proposes an algorithm to achieve the optimal tradeoff. Section V develops the ID-GAN algorithm. The performance of both methods is then assessed in Section VI, and Section VII concludes the paper.
@@ II Related Work
II-A Joint source coding and modulation (JSCM)
JSCM schemes outperform classical source-channel separation methods. Prior work on JSCM methods, also called JSCC even when they output complex-valued signals, can be divided into two categories according to the type of channel: basic channel transmission, in which the channel is simple like a Gaussian or Rayleigh channel, and advanced channel transmission, in which more realistic models for the channels are adopted, and the emphasis is on optimizing the transmission aspect of the system.
@@ II-A Joint source coding and modulation (JSCM)
JSCM for basic channels. Methods in this category typically focus on designing neural networks that optimize the compression performance of the JSCC/JSCM system, while neglecting aspects of transmission optimization, such as radio resource allocation. For example, [deep-jscc] proposed a deep joint source-channel coding (D-JSCC) algorithm based on an autoencoder and showed that besides outputting images with quality superior to separation-based schemes, the algorithm exhibits graceful performance degradation in low SNR. Techniques vary according to the domain of the data, e.g., text, image, video, or multimodal data. For example, the JSCM system designed in [Farsad18-DJSCC_Text] used a recurrent neural network for transmitting text. Also focusing on text transmission, [Xie21-DJSCC_Semantic] proposed a semantic communication system (DeepSC) based on a transformer and, to evaluate performance, also a novel metric to measure sentence similarity. DeepSC was extended in [Weng21-DJSCC_Speech] for speech transmission. JSCM has also been applied to the transmission of multimodal data. For instance, [Wan23-CooperativeMM] proposed a cooperative scheme to transmit audio, video, and sensor data from multiple end devices to a central server. And concentrating on text and images, [Zhang22-SemanticMM] designed a coarse-to-fine multitask semantic model using an attention mechanism. The theory and algorithms we derive in this paper fall under this category, as we consider Gaussian channels.
JSCM for advanced channels. Methods in this category adopt more realistic channel, like the erasure channel [DeepJSCC-l], feedback channel with channel state information (CSI) [DeepJSCC-f, Jialong23-DJSCC_CSI], and the waveform (OFDM, etc.) or multi-user channels. They focus on optimizing transmission. For example, [deep-edge] designed retrieval-oriented image compression schemes, [DeepJSCC-f] used channel feedback to improve the quality of transmission, and [DeepJSCC-l] considered an adaptive bandwidth to transmit information progressively under an erasure channel. Furthermore, [Jialong23-DJSCC_CSI] designed an end-to-end approach for D-JSCC [deep-jscc] with channel state information (CSI) feedback. The main idea was to apply a non-linear transform network to compress both the data and the CSI. Finally, [Yang21-DJSCC_OFDM] designed a scheme for orthogonal frequency division multiplexing (OFDM) transmission that directly maps the source images onto complex-valued baseband samples.
II-B GAN-based compression
@@
Most algorithms for image transmission are based on autoencoders [rumelhart1986learning], e.g., [deep-jscc, deep-edge, DeepJSCC-l, DeepJSCC-f, Jialong21-DSCC_Attention]. Autoencoders, however, compress signals only up to moderate compression ratios. Although they achieve high-quality reconstruction, this is at the cost of communication efficiency. Extreme compression has been achieved instead by using GANs [goodfellow2014generative], which are generative models that learn, without supervision, both a low-dimensional representation of the data and its distribution [karras2019style]. This gives them the potential to achieve extreme compression without undermining image perception quality. For example, [agustsson2019generative] proposed an autoencoder-GAN (AE+GAN) image compression system in which the encoder and decoder are trained simultaneously. The resulting method can achieve extremely low bitrates. One of the algorithms we propose, ID-GAN, requires less training (as encoder and decoder are trained separately), but attains a performance similar to or better than AE+GAN [agustsson2019generative].
@@
Also related to our work, [Erdemir23-Generative_JSCC] proposed two algorithms, inverse-JSCC and generative-JSCC, to reconstruct images passed through a fixed channel with a high compression ratio. The inverse-JSCC algorithm views image reconstruction as an inverse problem and uses a powerful GAN model, StyleGAN-2 [Karras20-StyleGAN2], as a regularizer together with a distortion loss that aligns with human perception, LPIPS [Zhang18-TheUnreasonableEffectivenessOfDeepFeatures]. It is thus an unsupervised method. Generative-JSCC transforms inverse-JSCC into a supervised method by learning the parameters of an encoder/decoder pair while keeping the parameters of StyleGAN-2 fixed. This work differs from ours in several ways. First, we consider not only distortion and perception metrics, but also classification accuracy and channel rate. In particular, the experiments in [Erdemir23-Generative_JSCC] do not consider any classification task. We also characterize the tradeoff between all these metrics. Second, our metric for perception, the Wasserstein-1 distance between the input and output distributions, differs from the LPIPS metric. Third, we train both the encoder and the decoder adversarially, while [Erdemir23-Generative_JSCC] uses a pre-trained GAN for the decoder. Finally, while training StyleGAN-2 in [Erdemir23-Generative_JSCC] (on a database of faces) requires tremendous computational resources, training our ID-GAN can be done with less resources.
II-C Tradeoff analyses